# Causal Representation Learning with Observational Grouping for CXR Classification

### v1

@@ 4 Discussion and Conclusion
In this work, we investigated grouping-based causal representation learning for disease classification in chest X-rays. By leveraging natural groupings in the data, our method improves the linear separability of disease features, leading to better predictive performance and lower error variability. We support these findings through latent space analysis and by observing consistent convergence behaviours across experiments. Our results, validated on multiple architectures and datasets, show that using our end-to-end invariant training strategy improves model stability and generalisation via identifiability. The choice of grouping attributes was guided by prior research on relevant demographic and non-demographic factors [5, 30]. Future work could explore applying this approach to other medical imaging tasks, multi-modal data, and alternative grouping criteria.
Acknowledgements
We thank Raghav Mehta, Fabio De Sousa Ribeiro and Pavithra Manoj for their detailed discussions and insightful feedback on early versions of this manuscript. R.R. is supported by the Engineering and Physical Sciences Research Council (EPSRC) through a Doctoral Training Partnerships PhD Scholarship. A.K. was supported by UKRI (grant no. EP/S023356/1), as part of the UKRI Centre for Doctoral Training in Safe and Trusted AI, and acknowledges support from the EPSRC Doctoral Prize. B.G. received support from the Royal Academy of Engineering as part of his Kheiron/RAEng Research Chair and acknowledges the support of the UKRI AI programme, and the EPSRC, for CHAI - EPSRC Causality in Healthcare AI Hub (grant no. EP/Y028856/1).

### v2

@@ 4 Discussion and Conclusion
In this work, we investigated grouping-based causal representation learning for disease classification in chest X-rays. By leveraging natural groupings in the data, our method improves the linear separability of disease features, leading to better predictive performance and lower error variability. We support these findings through latent space analysis and by observing consistent convergence behaviours across experiments. Our results, validated on multiple architectures and datasets, show that using our end-to-end invariant training strategy improves model stability and generalisation through identifiability. The choice of grouping attributes was guided by prior research on relevant demographic and non-demographic factors [gichoya2022ai, yang2024limits]. Future work could explore applying this approach to other medical imaging tasks (e.g. scanner invariance in breast imaging), multi-modal data (e.g. text-image pairs), and alternative grouping criteria (e.g. invariance to multiple groups based on assumptions from a causal graph). More broadly, grouping-based methods like ours can be interpreted as enforcing causal invariances to remove spurious shortcuts in classifiers [NEURIPS2022_d791394d]. Enforcing invariances may reduce accuracy or suppress meaningful group differences [zhao2022fundamental, petersen2023demographically], thereby compromising group-fairness. We therefore advise that how and when to use invariance strategies should be guided by domain-specific fairness analyses and expert knowledge of the causal data-generating process. Our code is available at https://github.com/RajatRasal/CRL-for-CXR-Classification.
{credits}
4.0.1 Acknowledgements
We thank Raghav Mehta, Fabio De Sousa Ribeiro, Pavithra Manoj and Fiona Kekwick for their detailed discussions and insightful feedback on early versions of this manuscript. R.R. is supported by the Engineering and Physical Sciences Research Council (EPSRC) through a Doctoral Training Partnerships PhD Scholarship. A.K. was supported by UKRI (grant no. EP/S023356/1), as part of the UKRI Centre for Doctoral Training in Safe and Trusted AI, and acknowledges support from the EPSRC Doctoral Prize. B.G. received support from the Royal Academy of Engineering as part of his Kheiron/RAEng Research Chair and acknowledges the support of the UKRI AI programme, and the EPSRC, for CHAI - EPSRC Causality in Healthcare AI Hub (grant no. EP/Y028856/1).
[…]
B.G. is a part-time employee of DeepHealth. No other competing interests.